scissor wrote:Hi Joe, my name is Greg.
I've just built my first LA500 and am experiencing a small problem. Like previous posters my comp works but appears to be suffering from the side chain oscillation problem. Voltage on audio out @ 1kHz is 1.23VAC. GR meter shows about -2dB (I'll let it warm up next time I re calibrate). If I have the input terminated with a muted mic pre, the comp ratio @ 3:1 and I wind up the threshold knob to about 2 o'clock the GR meter moves back to -20dB, then as I increase the threshold knob all the way clockwise the GR meter comes back to -10dB. This is only made worse with the two higher ratio settings and isn't affected at all by the side chain filter freq switch. During this process I can un mute the mic pre and monitor the audio going through the comp with headphones. Apart from the obvious attenuation, everything sounds fine.
As per your earlier advice, I've made sure the 3 wires to the threshold pot are twisted tightly and are as short as possible. I've been over both boards with a magnifying glass and cleaned them with Fluxclene.
How would you suggest I proceed from here?
First make sure you have C8 47pF fitted on the LA500 pcb under the JLM99v opamp.
Second make sure you have C2 100pF fitted on the MAC pcb.
If oscillation is still happening making the meter in GR mode show compression without anything plugged into the input of the LA500 when the threshold is turned up.
Add a 100pF cap at the threshold pot between the centre wiper leg and the 0v ground leg and all oscillation should stop.